# How Primordial Black Holes Change BBN

### v1

@@ 4 CHANGE IN BBN FROM PBH DECAY
While PBHs emit the full spectrum of Standard Model particles, not all of them propagate freely into the cosmological plasma. Whether an emitted particle significantly affects the background depends on its mean free path relative to the black hole’s characteristic scale. This criterion is key to identifying which emissions influence BBN. Previous studies [Vainer:1977, Vainer:1978] examined PBHs with masses in the range $10^{9}\,\mathrm{g}\leq M\leq 10^{12}\,\mathrm{g}$, and found that only photons, with mean free path $l_{\gamma}\sim 10^{-9}\,\mathrm{cm}$, and neutrinos, with $l_{\nu}\sim 10^{-11}\,\mathrm{cm}$ travel farther than the Schwarzschild radius of a PBH with $M\sim 10^{10}\,\mathrm{g}$ ($r_{\mathrm{s}}\sim 10^{-17}\,\mathrm{cm}$). The vicinity of a black hole would therefore be transparent only to photons and neutrinos, which can directly participate in background nuclear reactions (see Tab. 1 and 2).
For heavier particles, the mean free path is smaller than the size of the black hole. Previous works [Carr:1976zz, He:2022wwy] have therefore suggested that evaporating PBHs do not release such species as free-streaming particles. Instead, their emissions thermalize within an optically thick “hot envelope” surrounding the PBHs. High-energy particles undergo repeated scatterings with the background plasma, eventually reaching a new thermal equilibrium. Following this picture, we assume that massive particles influence BBN only indirectly: by depositing their energy into the surrounding medium, they heat the plasma and contribute to its entropy density.
@@
Once a viable combination $\beta(n_{s},\mathrm{d}n_{s}/\mathrm{d}\ln(k),M)$ is identified, it is incorporated into the BBN reaction network. In our framework, PBHs primarily act as sources emitting neutrinos and plasma particles through Hawking radiation. At high temperatures, before photons and neutrinos decouple (when neutrino–electron and neutrino–positron weak interactions remain efficient), all species share a common temperature governed by local thermal equilibrium. As the universe cools and photons and neutrinos decouple (at an equivalent of $T_{\mathrm{dec}}\approx 1.5\,\mathrm{MeV}$ in the absence of PBHs [Dodelson:2020bqr]), the neutrino and plasma sectors evolve separately. We make the assumption that neutrinos remain in Fermi–Dirac equilibrium with zero chemical potential, thereby neglecting energy down-scattering transport processes. Reference [Grohs:2020xxd] gives the prescription for evolving the temperature under such constraints. Meanwhile, the plasma particles continue to thermalize through Compton scattering, double Compton scattering, and bremsstrahlung. The non-neutrino energy density emitted by PBHs are assumed to instantaneously thermalize with the surrounding plasma, injecting entropy into the background, while the emitted neutrinos directly participate in background weak and nuclear reactions.
@@
Table 2: Nuclear reaction chain in the burst code [Grohs:2015eua, Grohs:2020xxd].
Because PBHs increase the comoving entropy density through Hawking radiation, we adjust the initial baryon-to-photon ratio $\eta$ at $T=30\,\mathrm{MeV}$ to a greater value than in standard BBN. This adjustment ensures that the final baryon-to-photon ratio at $T=8.6\times 10^{-4}\,\mathrm{keV}\approx 10^{4}\,\mathrm{K}$ matches the observed CMB value of $\eta=6.1\times 10^{-10}$ [Planck:2018vyg]. Practically, we determine a self-consistent $\eta$ by an iterative correction procedure analogous to that used for the scale factor $a(t)$ in Sec. 3: we evolve the system with an initial guess for $\eta$, compare the resulting value with the target, and repeat until the relative difference falls below $10^{-6}$. This yields consistent values of $a(t)$ and $\eta$ that fully incorporate both the PBH mass fraction and entropy injection from Hawking radiation.
@@ 5.1 Entropy and Primordial Abundances
Here we present the key findings of our calculations: the comoving entropy density increases as the universe expands and cools during the nucleosynthesis epoch. Previous studies constraining the PBH fraction using BBN generally neglected entropy variation, since they considered only extremely small PBH abundances ($\beta(M)\sim 10^{-20}$[Carr:2009jm, Acharya:2020jbv, Keith:2020jww], for which particle emission has a negligible effect on plasma heating. In contrast, our framework allows for ($\beta(M)\sim 1$), where PBH evaporation significantly alters the comoving entropy density, making it essential to account for this variation to ensure a self-consistent cosmological evolution in our PBH-BBN simulations.
@@
In each plot from Fig. 4 to 7, we observe that when the PBH mass fraction satisfies $\beta<10^{-10}$, the presence of PBHs has a negligible impact on the light element abundances because of their very small fraction. In this regime, the abundances converge to values consistent with a standard computation of BBN absent PBHs. Those predicted primordial abundances are consistent with the observed values, except for $X_{{}^{7}{\mathrm{Li}}}$, which remains roughly a factor of 5 higher than the observed value [Jedamzik:2007cp, Asplund:2005yt].
@@
The injection of high-energy photons through Hawking radiation heats the background photons, which are tightly coupled to baryons before decoupling. As a result, the average kinetic energy of all particle species in thermal equilibrium increases. The increased energy of charged particles raises the tunneling probability through the Coulomb barrier, further amplifying reaction rates. On the other hand, a hotter photon bath strengthens photodisintegration processes, which counteract nuclear synthesis. The larger net rates result in keeping the light-element abundances closer to nuclear statistical equilibrium [Burbidge:1957vc].
@@ 6 SUMMARY AND CONCLUSION
We again stress the key outcome of our analysis: PBH evaporation will substantially increase the comoving entropy density. Previous studies focused on relatively small PBH mass fractions ($\beta\sim 10^{-20}$) and neglected the plasma heating from Hawking radiation. In the presence of a larger PBH mass fraction, Fig. 3 clearly shows the magnitude and epoch of the entropy increase from PBH evaporation. Furthermore, the neutrino component of the Hawking radiation changes the neutron-to-proton interconversion rates as demonstrated in Figs. 911. Coupled together, the non-standard evolution of the entropy and neutrino densities induce the novel primordial abundance patterns visible in Figs. 47, none of which follow from simple scaling laws over the PBH mass range we study. Our findings highlight the necessity of incorporating entropy evolution self-consistently when modeling the gravitational and nuclear physics of PBH-BBN interactions.

### v2

@@ 4 CHANGE IN BBN FROM PBH DECAY
While PBHs emit the full spectrum of Standard Model particles, not all of them propagate freely into the cosmological plasma. Whether an emitted particle significantly affects the background depends on its mean free path relative to the black hole’s characteristic and inter-separation scale. This criterion is key to identifying which emissions influence BBN and may lead to inhomogeneous heating during BBN. Previous studies [74] examined PBHs with masses in the range $10^{9}\,\mathrm{g}\leq M\leq 10^{12}\,\mathrm{g}$, and found that only photons, with mean free path $l_{\gamma}\sim 10^{-9}\,\mathrm{cm}$, and neutrinos, with $l_{\nu}\sim 10^{-11}\,\mathrm{cm}$ travel farther than the Schwarzschild radius of a PBH with $M\sim 10^{10}\,\mathrm{g}$ ($r_{\mathrm{s}}\sim 10^{-17}\,\mathrm{cm}$). The vicinity of a black hole would therefore be transparent only to photons and neutrinos, which can directly participate in background nuclear reactions (see Tab. 1 and 2).
For heavier particles, the mean free path is much smaller, and therefore annihilate quickly around a cloud of PBH’s horizon size [74]. Within our approximation, ignoring inhomogeneous heating during BBN is justified. Previous works [17, 47] have therefore suggested that evaporating PBHs do not release such species as free-streaming particles. Instead, their emissions thermalize within an optically thick “hot envelope” surrounding the PBHs. High-energy particles undergo repeated scatterings with the background plasma, quickly reaching a new thermal equilibrium. A similar treatment of rapid thermalization of PBH emissions is used in [9]. Following this picture, we assume that massive particles influence BBN only indirectly: by depositing their energy into the surrounding medium, they heat the plasma and contribute to its entropy density. While we do not consider the direct interaction of injected hadrons/mesons with the background plasma, complementary analyses based on different assumptions instead track the hadronic/meson component explicitly and study its effect on the neutron–to-proton conversion rate [75].
@@
Once a viable combination $\beta(n_{s},\mathrm{d}n_{s}/\mathrm{d}\ln(k),M)$ is identified, it is incorporated into the BBN reaction network. In our framework, PBHs primarily act as sources emitting neutrinos and plasma particles through Hawking radiation. At high temperatures, before photons and neutrinos decouple (when neutrino–electron and neutrino–positron weak interactions remain efficient), all species share a common temperature governed by local thermal equilibrium. As the universe cools and photons and neutrinos decouple (at an equivalent of $T_{\mathrm{dec}}\approx 1.5\,\mathrm{MeV}$ in the absence of PBHs [30]), the neutrino and plasma sectors evolve separately. We make the assumption that neutrinos remain in Fermi–Dirac equilibrium with zero chemical potential, thereby neglecting energy down-scattering transport processes. Reference [38] gives the prescription for evolving the temperature under such constraints. Meanwhile, the plasma particles continue to thermalize through Compton scattering, double Compton scattering, and bremsstrahlung. The non-neutrino energy density emitted by PBHs are assumed to instantaneously thermalize with the surrounding plasma, injecting entropy into the background, while the emitted neutrinos directly participate in background weak and nuclear reactions.
@@
Table 2: Nuclear reaction chain in the burst code [37, 38].
Because PBHs increase the comoving entropy density through Hawking radiation, we adjust the initial baryon-to-photon ratio $\eta$ at $T=30\,\mathrm{MeV}$ to a greater value than in standard BBN. This adjustment ensures that the final baryon-to-photon ratio at $T=8.6\times 10^{-4}\,\mathrm{keV}\approx 10^{4}\,\mathrm{K}$ matches the observed CMB value of $\eta=6.1\times 10^{-10}$ [2]. Practically, we determine a self-consistent $\eta$ by an iterative correction procedure analogous to that used for the scale factor $a(t)$ in Sec. 3: we evolve the system with an initial guess for $\eta$, compare the resulting value with the target, and repeat until the relative difference falls below $10^{-6}$. This yields consistent values of $a(t)$ and $\eta$ that fully incorporate both the PBH mass fraction and entropy injection from Hawking radiation.
@@ 5.1 Entropy and Primordial Abundances
Here we present the key findings of our calculations: the comoving entropy density increases as the universe expands and cools during the nucleosynthesis epoch. Previous studies constraining the PBH fraction using BBN generally neglected entropy variation, since they considered only extremely small PBH abundances ($\beta(M)\sim 10^{-20}$[13, 1, 57], for which particle emission has a negligible effect on plasma heating. In contrast, our framework allows for ($\beta(M)\sim 1$), where PBH evaporation significantly alters the comoving entropy density, making it essential to account for this variation to ensure a self-consistent cosmological evolution in our PBH-BBN simulations.
@@
In each plot from Fig. 4 to 7, we observe that when the PBH mass fraction satisfies $\beta<10^{-10}$, the presence of PBHs has a negligible impact on the light element abundances because of their very small fraction. In this regime, the abundances converge to values consistent with a standard computation of BBN absent PBHs. Those predicted primordial abundances are consistent with the observed values, except for $X_{{}^{7}{\mathrm{Li}}}$, which remains roughly a factor of 5 higher than the observed value [52, 3].
@@
The injection of high-energy photons through Hawking radiation heats the background photons, which are tightly coupled to baryons before decoupling. As a result, the average kinetic energy of all particle species in thermal equilibrium increases. The increased energy of charged particles raises the tunneling probability through the Coulomb barrier, further amplifying reaction rates. On the other hand, a hotter photon bath strengthens photodisintegration processes, which counteract nuclear synthesis. The larger net rates result in keeping the light-element abundances closer to nuclear statistical equilibrium [10].
@@ 6 SUMMARY AND CONCLUSION
We again stress the key outcome of our analysis: PBH evaporation will substantially increase the comoving entropy density. Previous studies focused on relatively small PBH mass fractions ($\beta\sim 10^{-20}$) and neglected the plasma heating from Hawking radiation, while more recent analyses incorporate the entropy injection explicitly [9]. In the presence of a larger PBH mass fraction, Fig. 3 clearly shows the magnitude and epoch of the entropy increase from PBH evaporation. Furthermore, the neutrino component of the Hawking radiation changes the neutron-to-proton interconversion rates as demonstrated in Figs. 911. Coupled together, the non-standard evolution of the entropy and neutrino densities induce the novel primordial abundance patterns visible in Figs. 47, none of which follow from simple scaling laws over the PBH mass range we study. Our findings highlight the necessity of incorporating entropy evolution self-consistently when modeling the gravitational and nuclear physics of PBH-BBN interactions.